# ME-IGM: Individual-Global-Max in Maximum Entropy Multi-Agent Reinforcement Learning

### v3

@@ A B C
[…]
$\epsilon$8.5 -8.5 -8.5
[…]
$\epsilon$8.5 0.2 0.2
[…]
5 The Overall Framework: ME-IGM
Figure 3 illustrates the framework of ME-IGM, which consists of three types of networks. Agents 1 to $n$ share a network $f_{\theta}$ for predicting local Q-values. The Mixer network, which aggregates local Q-values and manages the credit assignment among agents, is parameterized through its associated hyper-network as $\text{Mixer}_{\theta}$ [^4] Our algorithm is compatible with any Mixer network design that satisfies the IGM condition. For instance, during evaluation, we utilize the Mixer networks of QMIX and QPLEX, leading to two variants of our algorithm: ME-QMIX and ME-QPLEX.. $[\text{OPT}_{i,\phi}]_{i=1}^{n}$ are $n$ order-preserving transformations, and $\text{OPT}_{i,\phi}(f_{\theta}(o_{t}^{i}),s_{t})$ defines a local stochastic policy for agent $i$. The update rules for these networks are described as follows.
@@ 5 The Overall Framework: ME-IGM
where $\tau_{t}=\{o_{t:t+1}^{1:n},u_{t:t+1}^{1:n},s_{t:t+1},r_{t}\}$, $Q_{tot_{\theta}}(\mathbf{u},s)=\text{Mixer}_{\theta}\big{(}f_{\theta}(o^{1},u^{1}),\ldots,f_{\theta}(o^{n},u^{n}),s\big{)}$, and the estimated Q-target $\widehat{\mathcal{T}}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u}_{t})=r_{t}+\gamma Q_{tot,\theta^{-}}(s_{t+1},\mathbf{u}_{t+1})$. As a common practice, we utilize a target Q network $Q_{tot,\theta^{-}}$ to stabilize training, with $\theta^{-}$ being the exponentially weighted moving average of $\theta$. Notably, local actions $u_{t:t+1}^{1:n}$ are sampled from local stochastic policies $\pi_{i}$ (as defined in Equation (10)) during roll-out, rather than selecting the argmax of local Q-functions. This offers two advantages: (1) the softmax policy function enhances exploration, and (2) incorporating state $s$ in the sampling process allows for better utilization of global information during centralized training.
Further, to train $[\text{OPT}_{i,\phi}]_{i=1}^{n}$, we adopt a sample-based version of Equation (7), denoted as $L(\phi)$:
@@
$\alpha_{\omega}$ is the temperature and, similar to Haarnoja et al. (2018), can be updated by minimizing the following loss function:
@@ 6 Experiments
Figure 4: The mean return and standard deviation of QMIX, QPLEX, ME-QMIX , ME-QPLEX in Overcooked. It shows that ME-QMIX and ME-QPLEX achieve higher returns and exhibit faster convergence, compared to QMIX and QPLEX which do not adopt maximum entropy.
@@
Protoss 5_vs_5 58.0 54.6 70.2 53.3 22.9 37.8 75.9$\pm$3.0 69.8$\pm$1.8 68.1 78.7
10_vs_10 58.3 58.0 69.0 53.7 12.1 8.6 78.5$\pm$2.7 78.7$\pm$3.4 59.6 79.8
10_vs_11 18.2 20.3 42.5 22.8 5.7 0.4 50.4$\pm$2.9 44.9$\pm$3.2 21.3 48.7
20_vs_20 38.1 44.5 69.7 27.2 - 2.0 73.9$\pm$3.0 38.3$\pm$4.0 76.3 80.6
20_vs_23 5.1 4.1 16.5 4.8 - 0.3 23.4$\pm$1.0 18.7$\pm$1.3 11.8 24.2
Terran 5_vs_5 52.0 56.2 58.4 70.0 37.9 57.1 70.2$\pm$2.6 70.0$\pm$4.8 53.3 69.9
10_vs_10 58.1 57.3 65.8 66.1 20.4 15.8 72.6$\pm$1.8 69.4$\pm$0.2 58.4 72.2
10_vs_11 28.6 31.0 39.4 41.4 9.1 7.5 49.6$\pm$1.6 51.0$\pm$0.4 28.4 53.9
20_vs_20 52.8 49.6 57.6 23.9 - 0.2 59.3$\pm$1.3 45.1$\pm$7.2 35.9 65.4
20_vs_23 11.2 10.0 10.0 7.0 - 0.0 18.7$\pm$0.7 17.4$\pm$2.9 4.7 17.7
Zerg 5_vs_5 41.0 37.2 37.2 47.8 29.1 35.2 50.8$\pm$4.6 52$\pm$3.9 48.6 55.0
10_vs_10 39.1 49.4 40.8 41.6 17.9 8.0 49.1$\pm$0.6 42.9$\pm$1.8 50.6 57.6
10_vs_11 31.2 26.0 28.0 31.1 14.0 1.5 32.6$\pm$1.7 45.3$\pm$0.6 34.8 41.5
20_vs_20 31.9 31.2 30.4 15.8 - 0.2 34.3$\pm$5.2 42.1$\pm$5.1 26.7 43.3
20_vs_23 15.8 8.3 10.1 6.7 - 0.2 14.8$\pm$6.4 17.5$\pm$3.2 8.2 21.3
Table 2: We record the average win rates of different algorithms on SMAC-v2. The results demonstrate that our method, ME-IGM, outperforms all previous baseline algorithms and achieves performance comparable to IL+MARL methods, despite not having access to expert opponent demonstrations or opponent modeling.
6.1 Evaluation on Matrix Games
@@
We first demonstrate the effectiveness of our algorithm in a classic one-step matrix game, shown as Table 1. It is a non-monotonic payoff matrix commonly used in previous studies (Son et al. 2019; Wang et al. 2020). We evaluate ME-QMIX, a variant of ME-IGM built upon QMIX, by comparing its performance against two state-of-the-art MARL methods: QMIX (Rashid et al. 2020), a value decomposition MARL method, and FOP (Zhang et al. 2021), a Maximum Entropy MARL method. We are concerned with whether the joint action with the highest selection probability corresponds to the globally maximal Q value, ensuring optimal returns when executing a deterministic policy.
@@
As shown in Table 1, QMIX suffers from high estimation error in payoff values of this non-monotonic game, due to its monotonicity constraints (as introduced in Section 2.3). Its deterministic policy converges to a suboptimal joint action (B, B) with a reward of 0. FOP, which employs a network architecture similar to QPLEX, is capable of accurately estimating value functions (i.e., the payoff). However, its policy tends to select suboptimal joint actions more frequently due to the misalignment between local policies and the maximum global Q-values, as discussed in Section 3. ME-QMIX exhibits smaller estimation errors for the payoff of optimal joint actions (which is our primary concern) while having larger errors for suboptimal ones. Crucially, the joint action with the maximum Q-value in ME-QMIX aligns with the true optimal joint action, and the local policy assigns the highest probability to selecting this optimal joint action, avoiding the misalignment issue through the use of OPTs.
6.2 Evaluation on Overcooked
@@
To demonstrate the advantage of employing the principle of maximum entropy, we compare ME-IGM with two leading value decomposition MARL methods (which do not involve maximum entropy): QMIX and QPLEX, on the Overcooked benchmark (Carroll et al. 2019) – a widely used coordination-focused multi-agent environment. In particular, we evaluate them in two representative layouts: cramped room and asymmetric advantages. These scenarios pose distinct challenges, including close collaboration under spatial constraints and asymmetric advantages in task division, and require coordinated exploration. For each scenario, we conduct repeated experiments with different random seeds to get the training curves of 10 million environment steps. As shown in Figure 4, ME-QMIX and ME-QPLEX exhibit significantly faster convergence and achieve higher performance levels in the early stages of training compared to their counterparts, QMIX and QPLEX.
6.3 Evaluation on SMAC-v2
@@
We further conduct a comprehensive evaluation of ME-IGM on the SMACv2 (Ellis et al. 2022) benchmark across 15 scenarios, comparing its performance against 8 baselines. The results, as shown in Table 2, are obtained by training each algorithm on each scenario for 101010 millions of environment steps. Each value in the table represents the average performance over three runs with different random seeds. ME-IGM outperforms the baseline algorithms across all scenarios by a significant margin. MAPPO and IPPO, as on-policy algorithms, have lower sampling efficiency, resulting in lower win rates within  [^10] M environment steps. QMIX and QPLEX, lacking exploration, tend to converge to local optimum. HASAC (Liu et al. 2023) and FOP perform poorly across all scenarios. Surprisingly, ME-IGM achieves performance on par with (or even surpassing) IL+MARL methods, despite not relying on expert opponent demonstrations or opponent modeling. The results of MAPPO, IPPO, QMIX, QPLEX, IMAX-PPO, InQ are taken from (Bui, Mai, and Nguyen 2024). A detailed comparison with actor-critic baselines is presented in the Appendix.
6.4 Ablation Study
Figure 5: Comparison of exploration strategies in CTDE MARL: ME-QMIX with maximum entropy vs. QMIX with epsilon-greedy and extended epsilon annealing period.
@@ 6.4 Ablation Study
Table 3: Our algorithm (+OPT) vs. its two ablated versions: +Entropy, obtained by removing OPTs, and QMIX, obtained by further removing softmax local policies.
@@
Figure 6: The impact of different hyperparameters for training $\alpha_{\omega}$ on the performance of ME-QMIX. (a) A lower learning rate for $\alpha_{\omega}$ corresponds to increased exploration. (b) ME-QMIX’s performance is not sensitive to the choice of the target entropy $\bar{\mathcal{H}}$ in Equation (10).
@@
Protoss 5v5 0.58±0.030.58$\pm$ 0.030.58 ± 0.03 0.74±0.040.74$\pm$ 0.040.74 ± 0.04
Terran 5v5 0.61±0.040.61$\pm$ 0.040.61 ± 0.04 0.70±0.050.70$\pm$ 0.050.70 ± 0.05
Zerg 5v5 0.41±0.040.41$\pm$ 0.040.41 ± 0.04 0.48±0.050.48$\pm$ 0.050.48 ± 0.05
Table 4: ME-QMIX (OPT) vs. its ablated version: ME-QMIX (MLP), where OPT is replaced with a 3-layer MLP.
@@ Acknowledgment
[…]
@@
Appendix B B. Proof of Theorem 4.1
Lemma B.1 (Joint Soft Policy Improvement).
@@
where $C=\underset{\mathbf{u}_{t+1}}{\max}|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))|.$ Due to the Pinsker’s inequality (Pinsker 1964), we have:
@@
Appendix C C. Hyper-network Structure
Figure 7: (a) The overall network architecture. Single Q-Net takes local observations as input and outputs a distribution over local actions $u^{i}$, also denoted as local $Q_{i}$. Functions OPTs are order-preserving transformations, ensuring the input and output dimensions are identical and have the same argmax values. Both the input and output of the OPTs are k-dimensional, where $k$ is the action dimension. The output of the OPT, after undergoing a softmax operation, becomes the policy, from which action $u_{t}^{i}$ is sampled. The Mixer network takes local $Q_{i}$ as input and outputs the global $Q_{tot}$. Green blocks utilize a recurrent neural network as the backbone, with all agents sharing parameters. For decentralized execution, only the green network is needed, and the maximum values of their output are selected. (b) The blue blocks’ network structure. The red blocks represent the hyper-network, which takes the global state $s_{t}$ as input and outputs the network’s weights and biases. The weights are constrained to be non-negative. The activation function is denoted by $\sigma(\cdot)$. The mixer network takes an n-dimensional Q function as input and outputs a one-dimensional $Q_{tot}$.

### v4

@@ A B C
$\epsilon$ 1-$2\epsilon$ $\epsilon$
$\epsilon$ -8.5 -8.5 -8.5
1-$2\epsilon$ -8.5 0.2 0.2
$\epsilon$ -8.5 0.2 0.2
5. The Overall Framework: ME-IGM
Figure 3 illustrates the framework of ME-IGM, which consists of three types of networks. Agents 1 to $n$ share a network $f_{\theta}$ for predicting local Q-values. The Mixer network, which aggregates local Q-values and manages the credit assignment among agents, is parameterized through its associated hyper-network as $\text{Mixer}_{\theta}$ [^3] Our algorithm is compatible with any Mixer network design that satisfies the IGM condition. For instance, during evaluation, we utilize the Mixer networks of QMIX and QPLEX, leading to two variants of our algorithm: ME-QMIX and ME-QPLEX.. $\text{OPT}_{i,\phi}]_{i=1}^{n}$ are $n$ order-preserving transformations, and $\text{OPT}_{i,\phi}(f_{\theta}(o_{t}^{i}),s_{t})$ defines a local stochastic policy for agent $i$. The update rules for these networks are described as follows.
@@ 5. The Overall Framework: ME-IGM
where the variables are defined as:
$\tau_{t}=\{o_{t:t+1}^{1:n},\;u_{t:t+1}^{1:n},\;s_{t:t+1},\;r_{t}\}$
$Q_{tot_{\theta}}(\mathbf{u},s)=\text{Mixer}_{\theta}\big(f_{\theta}(o^{1},u^{1}),\ldots,f_{\theta}(o^{n},u^{n}),s\big)$
and the estimated Q-target is:
$\widehat{\mathcal{T}}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u}_{t})=r_{t}+\gamma Q_{tot,\theta^{-}}(s_{t+1},\mathbf{u}_{t+1})$
As a common practice, we utilize a target Q network $Q_{tot,\theta^{-}}$ to stabilize training, with $\theta^{-}$ being the exponentially weighted moving average of $\theta$. Notably, local actions $u_{t:t+1}^{1:n}$ are sampled from local stochastic policies $\pi_{i}$ (as defined in Equation (10)) during roll-out, rather than selecting the argmax of local Q-functions. This offers two advantages: (1) the softmax policy function enhances exploration, and (2) incorporating state $s$ in the sampling process allows for better utilization of global information during centralized training.
Further, to train $\text{OPT}_{i,\phi}]_{i=1}^{n}$, we adopt a sample-based version of Equation (7), denoted as $L(\phi)$:
@@
$\alpha_{\omega}$ is the temperature and, similar to haarnoja2018soft, can be updated by minimizing the following loss function:
@@ 6. Experiments
Figure 4. The mean return and standard deviation of QMIX, QPLEX, ME-QMIX , ME-QPLEX in Overcooked. It shows that ME-QMIX and ME-QPLEX achieve higher returns and exhibit faster convergence, compared to QMIX and QPLEX which do not adopt maximum entropy.
[…]
The mean return and standard deviation of QMIX, QPLEX, ME-QMIX , ME-QPLEX in Overcooked.
Table 2. Average win rates of different algorithms on SMAC-v2. The results we recorded demonstrate that our method, ME-IGM, outperforms all previous baseline algorithms and achieves performance comparable to IL+MARL methods. Note that the results for MAPPO, IPPO, QMIX, QPLEX, IMAX-PPO, and InQ are cited directly from bui2024mimicking as mean values; therefore, no standard deviations are reported for these baselines.
@@
[…]
6.1. Evaluation on Matrix Games
@@
We first demonstrate the effectiveness of our algorithm in a classic one-step matrix game, shown as Table 1. It is a non-monotonic payoff matrix commonly used in previous studies son2019qtran; wang2020qplex. We evaluate ME-QMIX, a variant of ME-IGM built upon QMIX, by comparing its performance against two state-of-the-art MARL methods: QMIX (rashid2020monotonic), a value decomposition MARL method, and FOP (zhang2021fop), a Maximum Entropy MARL method. We are concerned with whether the joint action with the highest selection probability corresponds to the globally maximal Q value, ensuring optimal returns when executing a deterministic policy.
@@
As shown in Table 1, QMIX suffers from high estimation error in payoff values of this non-monotonic game, due to its monotonicity constraints (as introduced in Section 2.3). Its deterministic policy converges to a suboptimal joint action (B, B) with a reward of 0. FOP, which employs a network architecture similar to QPLEX, is capable of accurately estimating value functions (i.e., the payoff). However, its policy tends to select suboptimal joint actions more frequently due to the misalignment between local policies and the maximum global Q-values, as discussed in Section 3. This observation highlights that the misalignment between learned policies and optimal joint actions is not confined to IGM-based approaches. ME-QMIX exhibits smaller estimation errors for the payoff of optimal joint actions (which is our primary concern) while having larger errors for suboptimal ones. Crucially, the joint action with the maximum Q-value in ME-QMIX aligns with the true optimal joint action, and the local policy assigns the highest probability to selecting this optimal joint action, avoiding the misalignment issue through the use of OPTs.
6.2. Evaluation on Overcooked
@@
To demonstrate the advantage of employing the principle of maximum entropy, we compare ME-IGM with two leading value decomposition MARL methods (which do not involve maximum entropy): QMIX and QPLEX, on the Overcooked benchmark carroll2019utility – a widely used coordination-focused multi-agent environment. In particular, we evaluate them in two representative layouts: cramped room and asymmetric advantages. These scenarios pose distinct challenges, including close collaboration under spatial constraints and asymmetric advantages in task division, and require coordinated exploration. For each scenario, we conduct repeated experiments with different random seeds to get the training curves of 10 million environment steps. As shown in Figure 4, ME-QMIX and ME-QPLEX exhibit significantly faster convergence and achieve higher performance levels in the early stages of training compared to their counterparts, QMIX and QPLEX.
6.3. Evaluation on SMAC-v2
@@
We further conduct a comprehensive evaluation of ME-IGM on the SMACv2 ellis2022smacv2 benchmark across 15 scenarios, comparing its performance against 8 baselines. The results, as shown in Table 2, are obtained by training each algorithm on each scenario for $10$ millions of environment steps. Each value in the table represents the average performance over three runs with different random seeds. ME-IGM outperforms the baseline algorithms across all scenarios by a significant margin. MAPPO and IPPO, as on-policy algorithms, have lower sampling efficiency, resulting in lower win rates within 1010M environment steps. QMIX and QPLEX, lacking exploration, tend to converge to local optimum. HASAC (liu2023maximum) and FOP perform poorly across all scenarios. Surprisingly, ME-IGM achieves performance on par with (or even surpassing) IL+MARL methods, despite not relying on expert opponent demonstrations or opponent modeling. The results of MAPPO, IPPO, QMIX, QPLEX, IMAX-PPO, InQ are taken from bui2024mimicking. A detailed comparison with actor-critic baselines is presented in the Appendix.
6.4. Ablation Study
Figure 5. Ablation study on exploration strategies in QMIX. Comparison between ME-QMIX and a modified QMIX with extended epsilon annealing on the Protoss 5v5 map. The results show that simply increasing epsilon-greedy exploration is insufficient for improving performance. In contrast, ME-QMIX enables QMIX agents to perform more structured exploration through the maximum entropy framework, leading to higher returns.
[…]
Comparison of exploration strategies in CTDE MARL
@@ 6.4. Ablation Study
Table 3. Our algorithm (+OPT) vs. its two ablated versions: +Entropy, obtained by removing OPTs, and QMIX, obtained by further removing softmax local policies. Results on three SMAC-v2 scenarios show that removing OPTs (+Entropy) leads to lower win rates due to misalignment between local and global objectives, while incorporating OPTs ensures consistent and monotonic policy improvement.
@@
Figure 6. The impact of different hyperparameters for training $\alpha_{\omega}$ on the performance of ME-QMIX. (a) A lower learning rate for $\alpha_{\omega}$ corresponds to increased exploration. (b) ME-QMIX’s performance is not sensitive to the choice of the target entropy $\bar{\mathcal{H}}$ in Equation (10).
[…]
The impact of different hyperparameters on the performance of ME-QMIX
Table 4. Ablation study on the effect of the Order-Preserving Transformation (OPT). We compare ME-QMIX (OPT) with a variant that replaces OPT layers with a standard 3-layer MLP, which does not preserve the order of input values. Results on three SMAC-v2 scenarios show that ME-QMIX (OPT) consistently achieves higher win rates, highlighting that enforcing order preservation is essential for mitigating local–global misalignment and achieving stable policy improvement.
@@
Protoss 5v5 $0.58\pm 0.03$ $0.74\pm 0.04$
Terran 5v5 $0.61\pm 0.04$ $0.70\pm 0.05$
Zerg 5v5 $0.41\pm 0.04$ $0.48\pm 0.05$
@@ Acknowledgment
Appendix A TD($\lambda$)
In this work we use TD($\lambda$) to update the Q-value. In this paragraph, we will derive the expression of TD($\lambda$) in the context of Maximum Entropy Reinforcement Learning. Specifically, we aim to demonstrate the derivation of the following equation:
@@ Appendix A TD($\lambda$)
Appendix B Proof of Theorem 3
[…]
@@ Appendix B Proof of Theorem 3
where $C=\underset{\mathbf{u}_{t+1}}{\max}|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{jt}^{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))|.$ Due to the Pinsker’s inequality pinsker1964information, we have:
@@
B.1. Convergence Analysis
In this section, we provide a theoretical analysis of the convergence properties of the proposed ME-IGM algorithm. Our analysis leverages the theoretical framework of Regularized Modified Policy Iteration (RMPI) established in previous literature
B.1.1. Theoretical Foundation: Regularized MPI
As proved by Geist et al. (smirnova2019convergenceapproximateregularizedpolicy), the convergence of policy iteration with time-varying regularization can be established through the following theorem:
{theorem}
[RMPI Convergence] Consider the Regularized Modified Policy Iteration algorithm with time-varying regularization functions $\Omega_{t}$. Let the sequence $(\lambda_{t})_{t}$ be a uniform bound for $\Omega_{t}$, such that:
$\sup_{\pi}\|\Omega_{t}(\pi)\|_{\infty}:=\sup_{\pi,s}|\Omega_{t}(\pi(\cdot|s))|\leq\lambda_{t}.$ (26)
The gap between the utility of the learned policy after $N$ iterations, $V_{N,\Omega}$, and the optimal utility $V^{*}$ is bounded by:
$\|V_{N,\Omega}-V^{*}\|_{\infty}\leq\frac{2}{1-\gamma}\left(\Lambda_{N}+\gamma^{N}\|V_{0,\Omega}-V^{*}\|_{\infty}\right),$ (27)
where $\Lambda_{N}:=\left(1+\frac{1-\gamma^{m}}{1-\gamma}\right)\sum_{t=1}^{N-1}\gamma^{N-t}\lambda_{t}$. Notably, if $\lambda_{t}\to 0$ as $t\to\infty$, the algorithm converges to the optimal value function $V^{*}$.
B.1.2. Mapping ME-IGM to RMPI
The discrepancy induced by the Order-Preserving Transformation (OPT) and the decentralized policy update can be analyzed as a time-varying regularization term $\Omega_{t}(\pi)$ within the Regularized MPI framework (which we denote as $\epsilon$ in our derivation).
As derived in Theorem 3, the error term $\Omega_{t}(\pi)$ is bounded by:
$\lambda_{t}=\frac{\gamma}{1-\gamma}\mathbb{E}\left[\Delta_{Q}\cdot\sqrt{D_{KL}(\pi_{t-1}||\pi_{t})}\right].$ (28)
Here, $\Delta_{Q}$ represents the alignment gap between the global Q-value and the transformed individual Q-values: $\Delta_{Q}(s,u)\triangleq|Q_{tot}(s,u)-\sum_{i}OPT_{i}(Q_{i},s)|$. By substituting Eq. 28 into the expression for $\Lambda_{N}$ in Eq. 27, we can quantify the total error induced by the misalignment $\Delta_{Q}$.
B.1.3. Empirical Convergence Analysis
Our current empirical evidence strongly supports the convergence of ME-IGM:
  1. (1)
    Vanishing $\Delta_{Q}$: Empirical results show that $\Delta_{Q}^{2}$ decreases by three orders of magnitude during the training process, indicating that the OPT effectively aligns the decentralized policies with the global Q-function.
  2. (2)
    Stable Policy Updates: The term $D_{KL}(\pi_{t-1}||\pi_{t})$ remains non-increasing during training, suggesting that the policy updates become increasingly stable.
Given these observations, we can approximately view $\lambda_{t}\to 0$ as training progresses. Under this approximation, Theorem B.1.1 guarantees that the ME-IGM algorithm converges to the optimal value function $V^{*}$, fulfilling the IGM condition while enjoying the exploration benefits of maximum entropy.
Appendix C Hyper-network Structure
Figure 7. (a) The overall network architecture. Single Q-Net takes local observations as input and outputs a distribution over local actions $u^{i}$, also denoted as local $Q_{i}$. Functions OPTs are order-preserving transformations, ensuring the input and output dimensions are identical and have the same argmax values. Both the input and output of the OPTs are k-dimensional, where $k$ is the action dimension. The output of the OPT, after undergoing a softmax operation, becomes the policy, from which action $u_{t}^{i}$ is sampled. The Mixer network takes local $Q_{i}$ as input and outputs the global $Q_{tot}$. Green blocks utilize a recurrent neural network as the backbone, with all agents sharing parameters. For decentralized execution, only the green network is needed, and the maximum values of their output are selected. (b) The blue blocks’ network structure. The red blocks represent the hyper-network, which takes the global state $s_{t}$ as input and outputs the network’s weights and biases. The weights are constrained to be non-negative. The activation function is denoted by $\sigma(\cdot)$. The mixer network takes an n-dimensional Q function as input and outputs a one-dimensional $Q_{tot}$.